OpenStudy (anonymous):

Solve the Equation: 5/2y+6 + 1/y-2 = 3/y+3

OpenStudy (anonymous):

(5/(2y)) + 6 + (1/(y-2) = 3/(y+3) you mean?

OpenStudy (anonymous):

No, its a rational expression

OpenStudy (amistre64):

\[\frac{5}{2y+6} + \frac{1}{y-2} =\frac{3}{y+3}\] perhaps?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

ahh k than think carefully, you should make all the denominators equal :)

OpenStudy (anonymous):

amistre how you write demoniators and numerators like that o.O? i can't do that -.-'

OpenStudy (amistre64):

persistence

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

so how do I get the answer?

OpenStudy (amistre64):

\[\frac{5}{2y+6} + \frac{1}{y-2} -\frac{3}{y+3}=0\] http://www.wolframalpha.com/input/?i=5%2F%282y%2B6%29+%2B+1%2F%28y-2%29+-3%2F%28y%2B3%29%3D0

OpenStudy (amistre64):

there are a lot of ways to get the answer :)

OpenStudy (anonymous):

i wish i culd use calculator in algebra exams :P

OpenStudy (amistre64):

the main thing to do prolly is just drudge thru the multiplication and simplification ....

OpenStudy (amistre64):

clear the fractions one by one by multiplying by each denominator thruout in succession

OpenStudy (anonymous):

Im just rying to figure out how to work it out.. i know the answer.. im trying to help a friend.

OpenStudy (amistre64):

for example: \[\frac{2}{A}+\frac{3}{B}-\frac{4}{C}=0\] \[\frac{2(A)}{A}+\frac{3(A)}{B}-\frac{4(A)}{C}=(A)0\] \[2(B)+\frac{3A(B)}{B}-\frac{4A(B)}{C}=(B)0\] \[2B(C)+3A(C)-\frac{4AB(C)}{C}=(C)0\] \[2BC+3AC-4AB=0\]

OpenStudy (amistre64):

then solve for the "y" that is sticking in all the places

OpenStudy (anonymous):

blehhh ok i think that sorta helped